# Model-based Manipulation of Deformable Objects with Non-negligible Dynamics as Shape Regulation

### v1

@@ IV-A2 Test Equipment
A Franka Emika FR3 7-DoF manipulator arm was used to provide actuation of the object’s base in the experiments. When required, a Bota Systems SenseOne 6-axis force-torque sensor was assembled in between to measure the actuation applied. An Intel Realsense D345 camera was used to capture RGB images at 30Hz, from which the object state could be extracted. All hardware control and communication were managed through a PC running Ubuntu 20.04 and the ROS Noetic middleware. Figure 4 shows an overview of the test equipment setup.
Figure 4: The experimental setup used for validation of the proposed model and control architecture.
@@ IV-B Extraction of the Object State
Points on the object at $s=\{0,0.5,1\}$ and $d=0$ were identified with coloured markers, facilitating the detection of their pixel locations via HSV thresholding. Under the assumption that the object stays confined to the plane on which the manipulator endpoint moves, the 3D marker positions were determined by intersecting this plane with the line projected through the associated pixel coordinate.
@@
The floating base system configuration best matching the observed points on the object can be extracted by the numerical inverse kinematics algorithm 1, where $q_{\mathrm{O},init}$ is an initial guess, $p$ are the measured coordinates of the three marked points, and $h(q_{\mathrm{O}})$ and $J(q_{\mathrm{O}})^{+}$ evaluate the forward kinematics and Jacobian pseudoinverse respectively at the same locations. $\epsilon$ and $\Delta$ are simply tuned to a desired accuracy and convergence rate.
Algorithm 1 Numerical Inverse Kinematics
@@ IV-C Parameter Identification
The model contains several object-specific parameters, some of which are not easy to directly measure. Measurable parameters are the object length $L$, diameter $D$, total mass of the object body $m_{L}$, object base mass $m_{0}$ and endpoint mass $m_{1}$. Immeasurable parameters are the object stiffness $k$, object damping $\beta$, and curvature offset $\bar{\Theta}$. To determine the unknown parameters, an identification procedure was developed in two parts, separating out those that could be determined from the static case and those that require dynamic analysis.
@@
IV-D Model Validation Results
A polynomial of degree $n=1$ was used for the experimental validation: an affine curvature strain paramerisation [60] of the object shape, where $\Theta=(\theta_{0},\theta_{1})$. The parameter identification procedures laid out in Section IV-C were carried out using this model. The ultimately determined values are summarised in Table II, while the rest of this subsection will detail the practical considerations and intermediate results.
@@ IV-D Model Validation Results
IV-D1 Steady State
Static equilibrium datasets were collected for each of the objects at 23 orientations of $\phi$, angled at evenly spaced $\frac{\pi}{12}$ rad increments excluding $-\pi$. To minimize the influence of hysteresis on the object’s shape between measurements, they were returned to $\phi=0$ and shaken loose before moving to each measurement orientation. Configurations of $\Theta$ corresponding to the measured $p_{s}$, $p_{m}$ and $p_{e}$ were then extracted to use as input to the parameter identification formulation for $k$ and $\bar{\Theta}$ as described in Section IV-C. Note that the affine curvature model exhibits parameter-dependent multi-stability [61]. Here, it is generally safe to assume that the object will settle to an expected equilibrium if it begins in a nominal static state at $\phi=0$ and $\Theta$ close to $(0,0)$, and is rotated slowly within $-\pi\leq\phi<\pi$.
@@ IV-D1 Steady State
Following identification for the static case, we check how well the resulting model predicts the equilibrium behaviour of the real objects. Fig. 5a shows the results of this process for OB1, plotted for $\phi=\{0,\pm\frac{\pi}{4},\pm\frac{\pi}{2},\pm\frac{3\pi}{4}\}$ (7 of the 23 datapoints used). Here, the coloured crosses indicate the measured positions of $p_{m}$ and $p_{e}$, and the light orange curves represent the configurations of $\Theta$ from inverse kinematics, used in (18) to identify the parameters. The dark orange curves are the steady state of the model obtained by forward simulation using the identified parameters. A composite image of the actual object shapes at these angles is shown in Fig. 5b. The distance between the measured and modeled $p_{e}$ is plotted against the angle $\phi$ in Fig. 6 for the same object, providing a measure of the modeling accuracy.
@@
Figure 5: a) Comparison of OB1 shape with $\Theta$ extracted from measured equilibria (light orange), and from modeling after parameter identification with the same data (dark orange). b) Composite image showing several of the measured equilibria for OB1.
@@
Figure 6: Error between measured and modeled equilibrium endpoint at each angle $\phi$ in the steady state model validation for OB1.
@@
Equivalent figures for the other objects are not included for the sake of space, but are available here. The results show that the static equilibrium model provides quite an accurate estimate for the endpoint position in the range $\frac{-3\pi}{4}<\phi<\frac{3\pi}{4}$ for all object types, with an error below 0.05m in almost all cases. Up to $\phi=\pm\pi$, most types continue to only suffer relatively mild deterioration in accuracy, with some notable exceptions. Most egregiously, OB4 and OB6 display a sudden spike in error for $\phi<\frac{-3\pi}{4}$; on closer inspection, the model with these properties is no longer multi-stable at this point, with the sole equilibrium point biased to the right-hand side of the $xy$ plane due to the offset curvature. Another limitation of the model is highlighted by OB5, with its worst-case combination of length, stiffness, and endpoint mass resulting in high curvature near the base followed by straightening out along the length. The affine curvature model cannot describe this over a single segment.
IV-D2 Dynamic Response
@@
The data for dynamic identification was collected from a single evolution of a pendulum motion for each object with the base fixed. Fig. 7a shows a time lapse of the images captured for OB1, along with the detected positions of $p_{s}$, $p_{m}$ and $p_{e}$, as well as the $\Theta$ configurations from inverse kinematics projected into the camera frame. Note that this shows only the first swing, the full sets of data run until the oscillations cease.
This data was used to estimate the value of $\beta$, and the resulting fully identified model was then simulated from the same initial conditions. A plot comparing the evolution of $\Theta$ measured from the experiment, and generated by the simulation, is shown in Fig. 7b.
In Fig. 7c this simulated evolution is then projected back onto the same captured images to show more clearly how the simulated object shape compares to the real object in Cartesian space. Only every second frame is included for clarity, and it can be seen that after release the simulated object shape leads the actual motion of the object, which matches the phase difference visible in Fig. 7b.
@@
Figure 7: a) Time lapse of partial extracted dynamic evolution of $\Theta$ for OB1. b) Comparison of the measured and simulated evolutions of the $\Theta$ states. c) Simulated evolution of $\Theta$ projected over captured images.
Equivalent plots to Fig. 7b for each object are included here, generally showing similarly closely matching dynamic responses, though the simulation tends to lead the real data for most object types. The time lapse images were not created for each object, but videos are available here.
@@ V Experimental Validation of the Controller
The control framework was experimentally validated, focussing on control of the object endpoint as a practical goal specification.
@@ V-A Position
In the first case we look at positioning the object endpoint at a goal $p^{*}$ only, so that $\zeta=p_{e}$, and take the Euclidean distance error as the optimization cost. We use a circular constraint on the floating base coordinates to represent the workspace of the manipulator, specifically a circle of radius 0.5m centred at $(0,0.333)$ in $\{S_{B}\}$, which approximates the reach of the FR3. Since the modeling accuracy and planar assumption start to deteriorate at higher base orientations, it was also decided to limit the range of angles to $\frac{-3\pi}{4}\leq\phi\leq\frac{3\pi}{4}$.
The optimization problem is therefore:
@@
We used the controller to position the object endpoint by solving (20) at each of a grid of endpoint goals, with 0.1m spacing. The grid covers an area in $\{S_{B}\}$ of $-0.7\leq x\leq 0.7$ m and $0.05\leq y\leq 0.75$ m, although depending on the object type it may not be necessary to cover the full height in the $y$-direction once it exceeds the reachable range. The MATLAB fmincon solver was used to obtain the solutions, and results were obtained for OB1, OB2, OB3 and OB5, as their properties cover a wide range of high deformation behaviour. Additionally, a ’model-free’ reference was evaluated, where $\phi=0$ and the manipulator is positioned at a constant offset to $p^{*}$ (or as close as possible given the manipulator workspace constraint), obtained from the measured steady state $p_{e}$.
@@
Fig. 8 shows the manipulator solutions, and corresponding equilibrium configurations of the object model, used for one row of the workspace validation in both cases for OB1. Lines are drawn pairing the goals with the achieved modeled endpoints, and the red circle indicates the constraint placed on the manipulator endpoint position.
Figure 8: Solutions for one row of the endpoint positioning experiment using OB1, for the model-free reference case (top) and model-based case (bottom).
Based on these solutions, we assess the controller experimentally by measuring $p_{e}$ after moving the manipulator to each configuration, as demonstrated in the example image sequences in Figure 9. The results for OB1 are summarised in Figure 10a, where the grid colours correspond to the error between the goal $p^{*}$, located at the centre of each cell, and $p_{e}$. In the upper row of plots, the error is calculated to the modeled $p_{e}$ determined in the optimization, plotted as white dots; lines are drawn between these and the goals. In the lower row of plots, the error is calculated to the measured $p_{e}$, plotted as white dots; here the lines are drawn between the modeled and measured $p_{e}$.
@@
Figure 9: Example image sequences from the endpoint position control validation experiment, with overlay of modeled final object goal states. a) OB1, b) OB2, c) OB3, d) OB5.
Figure 10: a) Visualisation of endpoint error over the workspace for OB1. Error from goal to modeled endpoint (upper row) and error from goal to measured endpoint (lower row). Model-free reference case (left) and model-based case (right). b) Mean endpoint positioning error taken over the full workspace grid, summarised for all tested objects.
@@
Complete results for the four objects, including the solutions for each row and heatmap grid summaries are included here. For all objects it is clear that the endpoint reachability is significantly expanded compared to the reference case, as expected. The measured positioning accuracy in this increased workspace varies significantly, generally corresponding to decreased accuracy with the use of larger $|\phi|$. For further quantitative analysis, Fig. 10b plots the mean endpoint error over all measured grid points for each tested object, although it is noted that these values vary depending on the chosen area of grid coverage. Improvement in the model-based case is clearly demonstrated here, there is a reduction in the measured mean error compared to the reference case of 63.0%, 68.1%, 78.5% and 57.7% for the four objects respectively.

### v2

@@ IV-A2 Test Equipment
A Franka Emika FR3 7-DoF manipulator arm was used to provide actuation of the object’s base in the experiments. When required, a Bota Systems SenseOne 6-axis force-torque sensor was assembled in between to measure the actuation applied. An Intel Realsense D435 camera was used to capture RGB images at 30Hz, from which the object state could be extracted. All hardware control and communication were managed through a PC running Ubuntu 20.04 and the ROS Noetic middleware. Fig. 4 shows an overview of the test equipment setup.
Figure 4: The experimental setup used for validation of the proposed model and control architecture. Relevant components are highlighted. Note that for simplicity and reproducibility, we are connecting the object directly to the robot’s end effector instead of grasping it via a gripper.
@@ IV-B Extraction of the Object State
Points on the object at $s=\{0,0.5,1\}$ and $d=0$ were identified with coloured markers, facilitating the detection of their pixel locations via HSV thresholding. Under the assumption that the object stays confined to the plane on which the manipulator endpoint moves, the 3D marker positions were determined by intersecting this plane with the line projected through the associated pixel coordinate. The positions of the base, mid-point, and endpoint of the DLO are extracted for two purposes: to estimate the object’s curvature and to present results and validate our approach at different points along the DLO.
Figure 5: Distributions of computation times for the inverse kinematics algorithm 1, for various order of approximation. The assumed DoFs for the object are equal to the polynomial order plus one. The time taken to converge to a solution was calculated for 1000 values randomly selected throughout the $\theta$-space.
@@
The floating base system configuration best matching the observed points on the object can be extracted via the algorithm 1. The algorithm has the form of a CLIK (closed-loop inverse kinematics) strategy with an over-constrained task space, as the number of target outputs (measurements of the current shape) is higher than the configurations. There, $q_{\mathrm{O},init}$ is an initial guess, $p$ are the measured coordinates of the three marked points, and $h(q_{\mathrm{O}})$ and $J(q_{\mathrm{O}})^{+}$ evaluate the forward kinematics and Jacobian pseudoinverse respectively at the same locations. Note that the pseudoinverse is here to be intended as a minimum MSE solution. Finally, $\epsilon$ and $\Delta$ are tuned to a desired accuracy and convergence rate. The calculation time of the Inverse Kinematics algorithm, which was outlined in Algorithm 1 was tested using the parameters identified for OB1. The tests were conducted using polynomial models of order 0 to 4, with an initial guess of $\theta=0$. The input values of $p_{\mathrm{m}}$ and $p_{\mathrm{e}}$ were taken from the DLO and generated over the range $\pi\geq\theta_{i}\geq\pi$. The results of the benchmarked computation times for this algorithm are shown in Fig. 5. The tests were run on an Intel 12th Gen i7-12700H (20) @ 4.600GHz CPU. These results demonstrate that the calculation can feasibly be integrated into a real-time control system, with both accuracy and computation time adjustable by selecting the appropriate polynomial order to suit specific application requirements.
Algorithm 1 Iterative Estimation of the Shape (CLIK-like)
@@ IV-C Parameter Identification
The model contains several object-specific parameters, some of which are not easy to directly measure. Measurable parameters are the object length $L$, diameter $D$, total mass of the object body $m_{\mathrm{L}}$, object base mass $m_{0}$, and endpoint mass $m_{1}$. Immeasurable parameters are the object stiffness $k$, object damping $\beta$, and curvature offset $\bar{\Theta}$. To determine the unknown parameters, an identification procedure was developed in two parts, separating out those that could be determined from the static case and those that require dynamic analysis.
@@
IV-D Comparative analysis of polynomial curvature across different orders
In the proposed approach, the configuration of the slender object is parametrized using polynomial curvature functions (3). While this representation is exact in the infinite-dimensional case, it can be truncated to a polynomial of degree $n$ to yield a finite-dimensional object configuration suitable for real-world applications. In this section, we compare the use of constant, linear, and quadratic curvature models for describing the shape of a real object. For this comparative analysis, we use the object OB1 and identify the equivalent object parameters for the three cases. With a first-order curvature model, the object configuration can be uniquely identified from two points (Section IV-C). For higher-order polynomials, multiple solutions exist, but only one satisfies the equilibrium. We recover parameters by minimizing the error between measured and predicted marker positions over 23 configurations spanning base orientations in $\pi$, $\pi$], following strategies similar to [64].
The resulting parameters for the three cases are as follows: 1) $k=0.1726$ and $\bar{\Theta}=-0.0005$, 2) $k=0.197$ and $\bar{\Theta}=(0.142,1.24)$ and 3) $k=0.1695$ and $\bar{\Theta}=(0.1081,1.5017,0.0022)$. Fig. 6 presents the results obtained after parameter identification for the three cases. The lower part of Fig. 6 shows the errors at the middle point ($\Delta p_{m}$) and the end-effector of the object ($\Delta p_{e}$), while the upper part illustrates the object’s shape for some configurations. The average errors with the standard deviations at the middle point for the three cases are $11\pm 5.5$ cm, $3.5\pm 1.6$ cm and $1.2\pm 0.4$ cm, while at the endpoint they are $11\pm 6.7$ cm, $2.9\pm 1.5$ cm and $1.5\pm 0.5$ cm, respectively. These results effectively demonstrate that increasing the order of the polynomial significantly improves accuracy, particularly in capturing the object’s middle point. However, this improvement comes at the cost of increasing the polynomial order, which, in turn, increases the dimensions of the matrices describing the system.
Figure 6: Comparative analysis of polynomial curvature across different orders. The results are obtained using constant (purple line), linear (blue line), and quadratic curvature (orange line). In the upper part, we show the shapes for three different orientations with colored backgrounds, which correspond to the colored regions in the figure below.
IV-E Model Validation Results
A polynomial of degree $n=1$ was used for the experimental validation: an affine curvature strain parametrization [65] of the object shape, where $\Theta=(\theta_{0},\theta_{1})$. The parameter identification procedures laid out in Section IV-C were carried out using this model. The ultimately determined values are summarised in Table II, while the rest of this subsection will detail the practical considerations and intermediate results.
IV-E1 Steady State
Static equilibrium datasets were collected for each of the objects at 23 orientations of $\phi$, angled at evenly spaced $\frac{\pi}{12}$ rad increments excluding $\pi$. To minimize the influence of hysteresis on the object’s shape between measurements, they were returned to $\phi=0$ and shaken loose before moving to each measurement orientation. Configurations of $\Theta$ corresponding to the measured $p_{s}$, $p_{\mathrm{m}}$ and $p_{\mathrm{e}}$ were then extracted to use as input to the parameter identification formulation for $k$ and $\bar{\Theta}$ as described in Section IV-C. Note that the affine curvature model exhibits parameter-dependent multi-stability [66]. Here, it is generally safe to assume that the object will settle to an expected equilibrium if it begins in a nominal static state at $\phi=0$ and $\Theta$ close to $(0,0)$, and is rotated slowly within $\pi\leq\phi<\pi$.
@@ IV-E1 Steady State
Following identification for the static case, we check how well the resulting model predicts the equilibrium behaviour of the real objects. Fig. 7a shows the results of this process for OB1, plotted for $\phi=\{0,\pm\frac{\pi}{4},\pm\frac{\pi}{2},\pm\frac{3\pi}{4}\}$ (7 of the 23 datapoints used). Here, the coloured crosses indicate the measured positions of $p_{\mathrm{m}}$ and $p_{\mathrm{e}}$, and the light orange curves represent the configurations of $\Theta$ from inverse kinematics, used in (22) to identify the parameters. The dark orange curves are the steady state of the model obtained by forward simulation using the identified parameters. A composite image of the actual object shapes at these angles is shown in Fig. 7b. Equivalent figures for the other objects are not included for the sake of space, but are available in the extra appendix materials.
@@
Figure 7: Comparison of the measured and modeled object equilibrium shapes for OB1. a) Here the light orange curves represent the configurations with $\Theta$ extracted from the measurements (colored crosses), and the dark orange are those obtained from modeling after parameter identification. b) Composite image showing the corresponding real object shapes.
(a) OB1
(b) OB2
(c) OB3
(d) OB4
(e) OB5
(f) OB6
@@
Figure 8: Error between measured and modeled object endpoint $p_{\mathrm{e}}$ at equilibrium, when the object base is moved to each angle $\phi$ in the steady state model validation of Section IV-E1.
@@
The error between the measured and modeled object end effector position, $\Delta p_{\mathrm{e}}$, and relative error $\Delta p_{\mathrm{e}}/L$, are plotted against the angle $\phi$ in Fig. 8 for the each object, providing a measure of the modeling accuracy. The results show that the static equilibrium model provides quite an accurate estimate for the endpoint position in the range $\frac{-3\pi}{4}<\phi<\frac{3\pi}{4}$ for all object types, with an error below 0.05m (8%) in almost all cases. Up to $\phi=\pm\pi$, most types continue to only suffer relatively mild deterioration in accuracy, with some notable exceptions. Most egregiously, OB4 and OB6 display a sudden spike in error for $\phi<\frac{-3\pi}{4}$; on closer inspection, the model with these properties is no longer multi-stable at this point, with the sole equilibrium point biased to the right-hand side of the $xy$ plane due to the offset curvature. This result shows that the in-plane assumption holds for a large range of end effector angles. It also highlights the effect out-of-plane motion of the object has on positioning accuracy. Larger out-of-plane movement is seen at higher angles of $\phi$, due to the gravitational forces no longer keeping a sufficient proportion of the object mass below the grasp point, removing this stabilizing effect.
Another limitation of the model is highlighted by OB5, with its worst-case combination of length, stiffness, and endpoint mass resulting in high curvature near the base followed by straightening out along the length. The affine curvature model cannot accurately describe this over a single segment, as will be shown later.
IV-E2 Dynamic Response
@@
The data for dynamic identification was collected from a single evolution of a pendulum motion for each object with the base fixed. In these experiments, the robot was set to $\phi=0$, starting from a deformed position far from equilibrium to observe the evolution of the states of the object. Fig. 9a shows a time lapse of the images captured for OB1, along with the detected positions of $p_{s}$, $p_{\mathrm{m}}$ and $p_{\mathrm{e}}$, as well as the $\Theta$ configurations from inverse kinematics projected into the camera frame. Note that this shows only the first swing, the full sets of data run until the oscillations cease. This data was used to estimate the value of $\beta$, and the resulting fully identified model was then simulated from the same initial conditions. In Fig. 9b this simulated evolution is then projected back onto the same captured images to show how the simulated object shape compares to the real object in Cartesian space, with only every second frame included for clarity. The time lapse images were not created for each object, but videos are available with the extra materials.
@@
Figure 9: Partial time lapse comparison of the dynamic evolution of $\Theta$ for OB1. The configurations extracted from the images are shown in (a), and the simulated evolution, after using this data for identification, is projected over the captured images in (b), for every second frame only.
(a) OB1
(b) OB2
(c) OB3
(d) OB4
(e) OB5
(f) OB6
Figure 10: Comparison between $\Theta$ state evolutions undergoing pendulum motion from the dynamic model validation of Section IV-E2. The measured data in the graphs is extracted from the image frames, and the simulated traces are generated from the model using the same initial $\Theta$ and with $\dot{\Theta}(t=0)=0$.
Plots comparing the evolution of $\Theta$ measured from the experiment and generated by the simulation are shown in Fig. 10. Several of the objects demonstrate quite closely matching responses, though there are some discrepancies for others. This is most obvious when comparing some of the amplitudes, however it is worth noting that as the two components of $\Theta$ generally occur with opposite phase, these plots can be a misleading indication of how the object shapes in Cartesian space compare. The discrepancies in the steady-state results are attributed to the hysteresis effect of the DLO, which tends towards the offset configuration, $\bar{\Theta}$. The hysteresis effect causes slight alterations in this configuration producing the errors seen observed in the figure. The sensitivity of our model to the hysteresis effect is analyzed in a subsequent section. Conversely, the phase lead in these plots that tends to appear fairly minor for most objects, is more clearly visible in Fig. 9b. The phase lead results from a slight mismatch in the initial state guess used during the dynamic simulation, as observed in Fig. 10. This discrepancy is attributed to minor variations in the parameters of the DLO. These factors make it particularly challenging to establish an accurate initial guess for validating the model. To better evaluate the dynamic evolution of the system, we compare the time evolution of the mid-point and end-effector coordinates of the DLO as observed in the experiment with those generated by our model in simulation, as shown in Fig. 11.
IV-F Influence of hysteresis in our model
In this section, we have analyzed the effect of hysteresis for a specific object (OB1) to quantify this effect in our model. Our observations indicate that the hysteresis effect of DLO can be effectively captured by adjusting the offset $\bar{\Theta}$ in the elastic force term. In the model, this parameter was identified to minimize the error across various shapes and was kept constant. However, we observed that if this parameter $\bar{\Theta}$ is identified separately for each case, its value can vary in the range $2\bar{\theta}_{i}$,  $2\bar{\theta}_{i}$] considering that $\bar{\theta}_{i}$ is the nominal value (see Table II).
Figure 11: Comparison of the end-effector coordinates of the DLO to evaluate the dynamic evolution of the system for OB2.
Figure 12: Influence of hysteresis for different configurations within the observed experimental range. The nominal case is plotted with an orange line in the upper part, while the different shapes using hysteresis are shown in grey.
Taking this into account, we performed a sensitivity analysis to examine how the shapes and errors change within this range for different configurations. The results of this analysis are presented in Fig. 12 where we analyze the sensitivity of our model to the hysteresis effect under different orientations of the robot’s end-effector. This effect results in an average error of 5 $\%$ and 7.5 $\%$ of the DLO’s length at the midpoint and endpoint, respectively. The error can be significantly reduced by loosening the DLO through shaking prior to conducting an experiment.
@@ V Experimental Validation of the Controller
The control framework was experimentally validated, focusing on control of the object endpoint as a practical goal specification. Due to the underactuated nature of DLO, not all configurations are attainable, as they must satisfy the static equilibrium condition. Therefore, we propose solving the kinematic inversion of our model for a given desired task coordinate that satisfies the static equilibrium condition through an optimization process.
@@ V-A Position
In the first case we look at positioning the object endpoint at a goal $p^{*}$ only, so that $\zeta=p_{e}$, and take the Euclidean distance error as the optimization cost. We use a circular constraint on the floating base coordinates to represent the workspace of the manipulator, specifically a circle of radius 0.5m centred at $(0,0.333)$ in $\{S_{B}\}$, which approximates the reach of the FR3. Since the modeling accuracy and planar assumption start to deteriorate at higher base orientations, it was also decided to limit the range of angles to $\frac{-3\pi}{4}\leq\phi\leq\frac{3\pi}{4}$. The optimization problem is therefore:
@@
We used the controller to position the object endpoint by solving (24) at each point of a grid of endpoint goals, with 0.1m spacing. The grid covers an area in $\{S_{B}\}$ of $0.7\leq x\leq 0.7$ m and $0.05\leq y\leq 0.75$ m, although depending on the object type it may not be necessary to cover the full height in the $y$-direction once it exceeds the reachable range. The MATLAB fmincon solver was used to obtain the solutions, and results were obtained for OB1, OB2, OB3 and OB5, as their properties cover a wide range of high deformation behaviour. Additionally, a ’model-free’ reference was evaluated to compare our method with alternative strategy for positioning the DLO’s end-effector, without considering its curvature model. This strategy involves constraining $\phi=0$ and positioning the manipulator at a constant offset to $p^{*}$ (or as close as possible, given the manipulator’s workspace constraint), which is derived from the measured steady state $p_{e}$.
@@
Fig. 13 shows the manipulator solutions and corresponding equilibrium configurations of the modeled object for each row of the workspace validation in both cases for OB1. Lines are drawn pairing the goals with the achieved modeled endpoints, and the red circle indicates the constraint placed on the manipulator endpoint position.
Figure 13: Solutions generated by the optimization (24) for OB1 in the endpoint positioning experiment described in Section V-A. The model-free reference case is shown on the left, and the model-based case on the right. The endpoint goal positions are marked with small crosses, and lines are drawn to the achievable optimized solution.
Based on these solutions, we assess the controller experimentally by measuring $p_{e}$ after moving the manipulator to each configuration, as demonstrated in the example image sequences in Fig. 14.
The results of these experiments carried out throughout the whole grid are summarized in Fig. 15, where the grid colors correspond to the error between the goal $p^{*}$, located at the center of each cell, and $p_{e}$.
(a) OB1
(b) OB2
(c) OB3
(d) OB5
@@
Figure 14: Example image sequences from the endpoint position control validation experiment in Section V-A, showing the manipulator moving to the desired object configurations corresponding to target endpoint goal locations. The dashed overlays indicate the intended final object goal states generated by the model-constrained optimization (24).
@@
For all objects it is clear that the endpoint reachability is significantly expanded compared to the ’model-free’ reference case, as expected. The measured positioning accuracy in this increased workspace varies significantly, generally corresponding to decreased accuracy with the use of larger $|\phi|$. The model-based case is clearly an improvement based on these demonstrations. Measuring the mean endpoint error over all measured grid points, there is a reduction in the measured mean error compared to the ’model-free’ reference case of 63.0%, 68.1%, 78.5% and 57.7% for the four objects respectively.
(a) OB1
(b) OB2
(c) OB3
(d) OB5
Figure 15: Visualization of the endpoint positioning accuracy over the workspace for OB1, OB2, OB3 and OB5. The plots (a) - (d), show the results of experiments to compare the model-free strategy and our model-based controller with shape regulation. The centre of each cell was the goal of the controller $p^{*}$ (white star). The intensity of the cell color indicates the error for that goal $p_{e}$. The red dot shows the measured position of the objects’s endpoint. A white dashed line is drawn between each goal and the measured point
(a) OB4 top, OB6 bottom
(b) OB4 top, OB6 bottom
(c) OB4 top, OB6 bottom
Figure 16: (a) Solutions generated by the optimization (25) for the endpoint orientation control experiment in Section V-B, and composites of the corresponding real object states. The yellow cross indicates the intended endpoint position. (b) Error between the goal and measured object tip angles. (c) Position errors between the goal and measured object tip position.